# MIRAGE: Multimodal foundation model and benchmark for comprehensive retinal OCT image analysis

### v2

@@ 2.2 Segmentation datasets
This dataset for the segmentation of geographic atrophy (SGA) is a private dataset consisting of 965 samples consisting of OCT volumes as well as SLO and fundus autofluorescence (FAF) images. All samples come from 100 patients (184 eyes) diagnosed with GA who were part of a clinical study on natural GA progression conducted at the Medical University of Vienna, Austria [65]. The SLO and FAF images were acquired with a Spectralis device centered on the macula with a FOV of $6\times 6\leavevmode\nobreak\ \text{mm}^{2}$ and a resolution of $1024\times 1024$ pixels. OCT and SLO images were co-registered by the device, while FAF and SLO images were registered with an in-house image registration pipeline based on aligning retinal vessel segmentation [110]. All samples have GA en face masks annotated by a retinal expert on FAF images. In this work, we used only the SLO images and the corresponding GA masks. For privacy reasons, the dataset cannot be made publicly available.
@@ Appendix 5 Supplementary Note 5: Computational efficiency
Supplementary Table 15: Computational resources and data requirements for pretraining MIRAGE compared to other foundation models. The table shows the training time, the number of epochs, the batch size, the number of samples, the data type and the learning method (fully supervised [FS] or self-supervised [SS]) used for pretraining along with the used hardware.
@@
DINOv2 $>$41 days - 2048 142M RGB natural images SS 160x V100 (32GB)
MultiMAE - 1600 2048 1.28M RGB, Depth, Seg. masks SS 8x A100 (80GB)
@@
RETFound 14 days 800 1792 736k OCT SS 8x A100 (40GB)
MIRAGE 25 days 1600 256 261k OCT, SLO, Layers SS 1x A100 (80GB)

### v3

@@ 2.2 Segmentation datasets
This dataset for the segmentation of geographic atrophy (SGA) is a private dataset consisting of 965 samples consisting of OCT volumes as well as SLO and fundus autofluorescence (FAF) images. All samples come from 100 patients (184 eyes) diagnosed with GA who were part of a clinical study on natural GA progression conducted at the Medical University of Vienna, Austria [65]. The SLO and FAF images were acquired with a Spectralis device centered on the macula with a FOV of $6\times 6\penalty 10000\ \text{mm}^{2}$ and a resolution of $1024\times 1024$ pixels. OCT and SLO images were co-registered by the device, while FAF and SLO images were registered with an in-house image registration pipeline based on aligning retinal vessel segmentation [110]. All samples have GA en face masks annotated by a retinal expert on FAF images. In this work, we used only the SLO images and the corresponding GA masks. For privacy reasons, the dataset cannot be made publicly available.
@@ Appendix 5 Supplementary Note 5: Computational efficiency
Supplementary Table 15: Computational resources and data requirements for pretraining MIRAGE compared to other foundation models. The table shows the training time, the number of epochs, the batch size, the number of samples, the data type and the learning method (fully supervised [FS] or self-supervised [$S$]) used for pretraining along with the used hardware.
@@
DINOv2 >>41 days - 2048 142M RGB natural images $S$ 160x V100 (32GB)
MultiMAE - 1600 2048 1.28M RGB, Depth, Seg. masks $S$ 8x A100 (80GB)
@@
RETFound 14 days 800 1792 736k OCT $S$ 8x A100 (40GB)
MIRAGE 25 days 1600 256 261k OCT, SLO, Layers $S$ 1x A100 (80GB)